# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v1

@@ 5.1 Task or domain ID
Most of the approaches specifically require information about task/domain ID, i.e., which domain/task the current sample belongs to. This information is then used to select or activate the task-specific branch in the model for classification/segmentation/feature extraction. If task/domain ID is needed then the approach is regarded as (a) task/domain aware else (b) task/domain agnostic. However, a well-designed CL approach should not rely on a task/domain ID to perform prediction Aljundi et al. (2017); Bayasi et al. (2021). Most of the approaches presume a-priori knowledge of domain IDs Derakhshani et al. (2022); Zhang et al. (2023b) while some infer from the data Zhang et al. (2023c); Bayasi et al. (2021); Aljundi et al. (2017). A very few do not require this information Derakhshani et al. (2022); González et al. (2022); Perkonigg et al. (2022).
@@ 5.2 Task boundaries
Typically, a shift is regarded as a change of data source, i.e., datasets coming from different sites or acquisition environments. Therefore, there are well-separated domain shifts or task boundaries. Thus, the necessary measures to adapt to the changed domain are triggered when there is a change in task/domain. Most of the works in the literature follow the assumption of well-separated task/domain boundaries.
However, in real-life cases, the change in domain/task could be interleaved or smooth rather than an abrupt transition. Such settings come under the umbrella of ‘blurred boundary continual learning’ Wang et al. (2023a). To better visualize the rigid and non-rigid boundary-based sequence of domains, a demonstration is provided via Figure 7. Recently a few works Hofmanninger et al. (2020a); Srivastava et al. (2021); González et al. (2022); Perkonigg et al. (2022) have explored such settings. For example González et al. (2022) consider two kinds of non-rigid shift in hippocampus segmentation datasets. In the first kind, termed ‘shifting source’, they create slowly shifting distributions for the segmentation task with three datasets (HarP, Dryad, and Decathlon) by interleaving samples from these datasets. In another case termed ‘transformed’, they create shift using TorchIO library Pérez-García et al. (2021) which performs various intensity rescaling and affine transform in the Decathlon dataset to create 5 episodes. Another example of a non-rigid boundaries-based CL scenario is given by Srivastava et al. (2021) for the chest x-ray classification problem. They consider three multi-site datasets (NIH Chest-X-rays14, PadChest, and CheXpert) as three domains; however, they curate a smooth transition from one dataset to another at the boundaries.
@@ 5.3 Sample annotation
Apart from information regarding the current task, annotation for the samples in the current task is also a major concern. Almost all the available CL strategies in the medical domain require sample annotation for learning the classification or segmentation task. It is undeniable that labeling is highly costly in the medical domain, especially in the histopathology field, and hence may hinder the rapid advancement of the field. In other domains, unsupervised CL strategies have been explored for example Ashfahani and Pratama (2022) only use a few labeled samples to associate clusters to classes and the learning is completely unsupervised manner.
@@ 6 Evaluation Strategy and Metrics
To evaluate a CL strategy various aspects like the performance of the model on past, current, or future episodes, resource consumption Zhang et al. (2023b), memory size Hofmanninger et al. (2020a), model size growth, execution time, etc can be explored. CL frameworks involving dynamic architecture or memory also report time and memory analysis. For example González et al. (2022) maintains model history as mean and variance parameters of multivariate Gaussian which also dynamically grow in the presence of drift; hence authors also provided training time analysis. If the framework does not involve dynamic architecture, mainly performance on data is reported. Further, the performance is mainly reported for evaluating episode performance, stability, and plasticity.
In CL, there is a given sequence of episodes ($1,...,T$) to sequentially learn. Typically we have well-defined training and test pairs as $(Tr_{i},Te_{i})$ for each $i^{th}$ episode. Upon sequential training, a train-test performance matrix $P\in\mathit{P}^{T\times T}$ is generated for the given sequence of episodes ($1,...,T$). An example matrix for $T=4$ is shown in Figure 8. The cell $P_{t,i}$ is performance on test data of $i^{th}$ episode when the training up-to training data of $t^{th}$ episodes is complete. Thus, after training up to $t^{th}$ episodes, the performance $P_{t,i}$ for $t<i$, $t=i$, and $t>i$ correspond to past, current, and future performance, respectively.
The performance metric is chosen depending on the application. For example, IoU, dice score/coefficient, SSIM, Hausdorff distance, and average symmetric surface distance, etc are popular choices for segmentation whereas accuracy, recall, F1-score, and AUC are usually computed for classification applications. Once we have chosen our performance metric, we can compute metrics for measuring stability and plasticity on top of it. Various metrics have been proposed to quantify stability (Table 11) and plasticity (Table 12); however all the metrics are derived from the same train-test performance matrix (Figure 8) to measure the amount of forgetting or forward transfer.
@@ 6.1 Stability: Backward Transfer
It refers to the catastrophic forgetting of a model upon learning new episodes. A model is regarded as stable if it gives a similar performance for a particular episode. However, upon learning the new episode naively, it is possible to forget the past learning and hence disrupt the performance of the previous episode. Whereas a carefully designed CL strategy will cause less forgetting or improvement on past episodes upon learning the new episodes. A backward transfer (BWT) metric is used to measure the amount of forgetting. In other words, we can say it is a way to measure stability, i.e., how well the model would retain the previously acquired knowledge to prevent catastrophic forgetting.
Lopez-Paz and Ranzato (2017) define BWT as the influence of new learning $t_{i}$ task on a previous task. When learning a task $t_{i}$ increases the performance on a future task, it is $+ve$ BWT and if it causes deterioration, it results in $-ve$ BWT. Large $-ve$ BWT is also termed as catastrophic forgetting. The BWT metric given by Lopez-Paz and Ranzato (2017) is very frequently adopted Srivastava et al. (2021); Hofmanninger et al. (2020a) and it is computed after finishing all the episodes as shown in Eq.(A) in Table 11. Average forgetting Derakhshani et al. (2021) is another popularly used metric Derakhshani et al. (2022); Chen and Tang (2022). It is computed as the average of the difference in highest performance and the final performance reached after training on all episodes is finished (Eq.(B) in Table 11). Chee et al. (2023) compute BWT after learning each subsequent task ($t^{th}$) rather than computing it at the last episode (Eq.(B) in Table 11) as shown through Eq.(C) in Table 11. González et al. (2022) compute BWT as the change in performance after training with each subsequent task averaged over the number of tasks (Eq.(D) in Table 11). Note that BWT is not defined for the last task. Another BWT metric contribution is by Özgün et al. (2020) as shown via Eq.(E) in Table 11.
@@ 6.2 Plasticity: Forward Transfer
The plasticity of a model is reflected in its capability to accommodate more and more knowledge or exploit the knowledge learnt so far to learn the new. In literature, this is frequently referred to as forward transfer (FWT). It is also said that FWT measures the “zero-shot" learning capability of a model Lopez-Paz and Ranzato (2017). In contrast to the usage of BWT, not all works report FWT. Certain approaches might enforce rigorous regularization on the model, which can result in poor performance on new tasks, consequently causing a comparatively reduced level of forgetting, i.e., high $+ve$ BWT. Therefore reporting FWT along with BWT is equally important for unbiased evaluation of the model.
The very popular FWT metric Srivastava et al. (2021); Hofmanninger et al. (2020a) shown via Eq.(F) in Table 12 is proposed by Lopez-Paz and Ranzato (2017) that quantifies FWT as the impact of learning a task $t_{i}$ on a future task. It is the average of the difference in the performance on $t_{i}$ task before learning it and the performance given by a model with random weights ($\bar{b}_{i}$). Note that FWT is not defined for the first task. Further, González et al. (2022) computes FWT as the change in performance in each stage before and upto $Tr_{i}$, averaged over the number of tasks (Eq.(G) in Table 12). Note that it is not defined for the first task. Another metric for FWT was proposed by Shu et al. (2022) as shown via Eq.(H) in Table 12. Özgün et al. (2020) defines two separate forward transfer metrics one for unseen (average of elements above diagonals of $P$) and seen episodes shown in Eq.(I) and Eq.(J) in Table 12, respectively. $b_{i}$ is the performance on a standalone model only trained on $i^{th}$ episode. Further, they report a “transfer-learning metric" as the sum of diagonal elements in matrix $P$ as measures of plasticity, i.e., the ability to adapt to new tasks.
@@ 6.3 Episode performance and others
Thus, after processing all the episodes, we can get total $T$ average accuracy values which we can plot and visualize for different approaches Derakhshani et al. (2022); Kaustaban et al. (2022). Further, researchers Derakhshani et al. (2022) also compute mean over these values ($\frac{1}{T}\sum_{i=1}^{T}[\frac{1}{t}\sum_{i=1}^{t}P_{t,i}$]) to get a single scale value which can be used to get a single scalar value to directly compare with other approaches. Another popular definition of average accuracy is to compute the above-mentioned metric only after finishing all the episodes Lopez-Paz and Ranzato (2017); Chen and Tang (2022); Kaustaban et al. (2022) as below.
@@
Ravishankar et al. (2019)
Ultrasound, X-ray
Cardiac view, Pneumothorax
DIS, TIS, data-incremental
Rehearsal + architectural
CNN
Supervised
Accuracy, BWT
Subset of ChestXRay (Wang et al. 2017)
Hofmanninger et al. (2020a)
CT
Chest
DIS
Rehearsal
Res-Net50 (He et al. 2016)
Supervised
Accuracy, BWT, FWT
Private
Lenga et al. (2020)
X-ray
Chest
DIS
Comparitive: EWC, LWF
DenseNet121 (Huang et al. 2017)
Supervised
AUC, BWT, FWT
ChestX-ray14, MIMIC-CXR
Li et al. (2020)
Dermoscopic
Skin disease
CIS
Dual-distillation+rehearsal
ResNet18, AlexNet, VGG19
Supervised
Accuracy
Skin8, Skin40 from ISIC2019 (Tschandl et al. 2018), CIFAR100
Srivastava et al. (2021)
X-ray
Chest
DIS
Rehearsal
ResNet101
Supervised
AUC, FWT, BWT
@@
NIH Chest-X-rays14 (McDermott et al. 2020), PadChest (Bustos et al. 2020), CheXpert (Wang et al. 2017)
Bayasi et al. (2021)
Dermoscopic
Skin lesion
DIS
Architectural (fixed but partitioned network)
ResNet-50
Supervised
Accuracy
HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
Morgado et al. (2021)
Dermoscopic
Dermatological imaging modality
DIS
Regularization, rehearsal
VGG-16, MobileNetV2
Supervised
Accuracy, precision, recall, F1-score, FWT, BWT
Private
Shu et al. (2022)
Fundus images
Eyes
DIS
Rehearsal + regularization
Resnet-18, 2-layer MLP (Saha et al. 2021)
Supervised
F1-score, BWT, FWT
@@
ODIR, RIADD, REFUGE, iSee (Fang et al. 2020), EyeQ (Fu et al. 2019), PMNIST (Rupesh Kumar et al. 2013)
Akundi and Sivaswamy (2022)
X-ray
Chest
CIS
Regularization
DenseNet121
Supervised
AUC, BWT
CheXpert
Kaustaban et al. (2022)
Histopathology
Tumor (breast cancer, colorectal cancer)
Data incremental, DIS, CIS, TIS
Regularization, rehearsal (comparative :EWC, LwF, AGEM, CoPE, iCaRL)
ResNet-18
Supervised
Accuracy, FWT, BWT
CRC (Kather et al. 2019), PatchCam (Bejnordi et al. 2017)
Chen and Tang (2022)
Histopathology
Breast cancer
CIS
Regularization
Alexnet, Densenet201, Resnet152
Supervised
Accuracy, forgetting
BreakHis
Derakhshani et al. (2022)
Multiple
Multiple disease
TIS, CIS, cross-DIS
Regularization, rehearsal (comparative: EWC, MAS, LwF, iCaRL, EEIL)
ResNet18
Supervised
Accuracy, forgetting
MedMNIST (Yang et al. 2023a)
Chee et al. (2023)
Histopathology, retinal, skin lesions
Colorectal cancer, eye, skin
CIS
Rehearsal + architecture + regularization
ResNet18
Supervised
Accuracy, forgetting
CCH5000 (Kather et al. 2016), EyePACS (kag ), HAM10000 (Tschandl et al. 2018)
Zhang et al. (2023c)
multiple
Disease (multiple organ source)
CIS
Rehearsal + architectural
ResNet18
Supervised
MCR
Skin8, Path16, CIFAR100
Bándi et al. (2023)
Histopathology
Cancer (multi-organ: breast, colon, head-neck)
DIS
Comparative: (Architecture: PackNet) (Regularization: EWC, GEM)
DenseNet
Supervised
FROC, ROC, Cohen’s Kappa
CAMELYON16, CAMELYON17, private for colon and head-neck
Bai et al. (2023)
VQA
Surgical visual-question localized-answering
CIS
Regularization
VisualBERT
Supervised
Accuracy, IoU
EndoVis18, EndoVis17, M2CAI
Sadafi et al. (2023)
microscopic
WBC
DIS, CIS, DIS + CIS
Rehearsal + regularization
ResNeXt-50
Supervised
Accuracy
Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Yang et al. (2023b)
Multiple
Disease classification
DIS+CIS
Rehearsal
ResNet18
Few-shot
Accuracy, performance dropping rate
MedMNIST (Yang et al. 2023a) (PathMNIST, DermaMNIST, OrganAMNIST, RetinaMNIST, BreastMNIST, BloodMNIST)
@@
Karani et al. (2018)
MRI
Brain
DIS
Architecture
U-Net
Supervised
DSC
@@
HCP (Van Essen et al. 2013), ADNI (ADNT ), ABIDE (Di Martino et al. 2014) and IXI (IXI )
Baweja et al. (2018)
MRI
@@
Normal brain structures, white matter lesions
TIS
Regularization: EWC
DeepMedic 3D convnet (Kamnitsas et al. 2017)
Supervised
DSC coefficient
UK Biobank (Miller et al. 2016)
McClure et al. (2018)
SMRI
Axial and sagittal brain
DIS
Regularization: DWC
MeshNet (Fedorov et al. 2017) (CNN-based)
Supervised
DSC
HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017)
Ozdemir et al. (2018)
MRI
Humerus and scapula
CIS
Regularization (LwF) + rehearsal
U-Net
Supervised
DSC, SurfDist
Private
Ozdemir and Goksel (2019)
MRI
Knee
CIS
Regularization
U-Net
Supervised
DSC, MSD
@@
SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
van Garderen et al. (2019)
4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR
Brain tumor
DIS
Regularization: EWC
3D U-Net CNN
Supervised
Dice score
BraTS 2018 (Bakas et al. 2017; Menze et al. 2014), private
Özgün et al. (2020)
MRI
Brain
DIS
Regularization: MAS
QuickNAT (Roy et al. 2019)
Supervised
CL-DSC, REM, BWT+, TL, FWT
@@ 4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR
CANDI (Kennedy et al. 2012), MALC (Asman and Landman 2013), and ADNI (Jack Jr et al. 2008)
Memmel et al. (2021)
MRI
Hippocampal
DIS
U-Net
Supervised
IoU, DSC
(Simpson et al. ), (Kulaga-Yoskovitz et al. 2015), (Boccardi et al. 2015)
Perkonigg et al. (2021)
MRI
Brain
DIS
Rehearsal
3D-ModelGenesis (Zhou et al. 2021)
Supervised
MAE, FWT, BWT
IXI (IXI ), OASIS3 (LaMontagne et al. 2019)
Ranem et al. (2022)
MRI
Hippocampus
DIS
Comparative (Chaudhry et al. 2018b; Cermelli et al. 2020; Douillard et al. 2021, 2020)
ViT U-Net
Supervised
DSC, FWT, BWT
@@
Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
Liu et al. (2022)
CT
@@
Liver, spleen, pancreas, right kidney left kidney
CIS
Rehearsal + regularization
nnUNet
Supervised
DC, HD95
(Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
Karthik et al. (2022)
MRI
@@
Multiple sclerosis lesions (brain)
DIS
Rehearsal
3D U-Net
Supervised
DSC, BWT
(Kerbrat et al. 2020)
González et al. (2022)
MRI
Hippocampus
DIS
Architectural
nnUNet (Isensee et al. 2021)
Supervised
DSC, FWT, BWT
@@
HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019)
Li et al. (2022)
CMR
Cardiac
DIS
Rehearsal
U-Net
Supervised
DSC, HD95, BWT, FWT
M&Ms (Campello et al. 2021)
Zhang et al.; Zhang et al. (2023d; 2021)
MRI
Prostate, optic cup and disc
DIS
Regularization
U-Net
Supervised
DSC, ASD

### v3

@@ 5.1 Task or domain ID
Most of the approaches specifically require information about task/domain ID, i.e., which task/domain the current samples belong to. This information is then used to select or activate the task-specific branch in the model for classification/segmentation/feature extraction. If task/domain ID is needed then the approach is regarded as (a) task/domain aware else (b) task/domain agnostic. However, a well-designed CL approach should not rely on a task/domain ID to perform prediction Aljundi et al. (2017); Bayasi et al. (2021). Most of the approaches presume a-priori knowledge of domain IDs Derakhshani et al. (2022); Zhang et al. (2023b) while some infer from the data Zhang et al. (2023c); Bayasi et al. (2021); Aljundi et al. (2017). A very few do not require this information Derakhshani et al. (2022); González et al. (2022); Perkonigg et al. (2022).
@@ 5.2 Task boundaries
Typically, a shift is regarded as a change of data source, i.e., datasets coming from different sites, acquisition environments, etc. Therefore, there are well-separated domain shifts or task boundaries. Thus, the necessary measures to adapt to the changed domain are triggered when there is a change in task/domain. Most of the works in the literature follow the assumption of well-separated task/domain boundaries.
However, in real-life cases, the change in domain/task could be interleaved or smooth rather than an abrupt transition. Such settings come under the umbrella of ‘blurred boundary continual learning’ Wang et al. (2023a). To better visualize the rigid and non-rigid boundary-based sequence of domains, a demonstration is provided via Figure 7. Recently a few works Hofmanninger et al. (2020a); Srivastava et al. (2021); González et al. (2022); Perkonigg et al. (2022) have explored such settings. For example, González et al. (2022) considers two kinds of non-rigid shifts in hippocampus segmentation datasets. In the first kind, termed ‘shifting source’, they create slowly shifting distributions with 3 datasets (HarP, Dryad, and Decathlon) by interleaving samples from these datasets for segmentation application. In another case termed ‘transformed’, they create shift using TorchIO library Pérez-García et al. (2021) which performs various intensity rescaling and affine transform in the Decathlon dataset to create five episodes. Another example of a non-rigid boundaries-based CL scenario is given by Srivastava et al. (2021) for the chest x-ray classification problem. They consider 3 multi-site datasets (NIH Chest-X-rays14, PadChest, and CheXpert) as 3 domains; however, they curate a smooth transition from one dataset to another at the boundaries.
@@ 5.3 Sample annotation
Apart from information regarding the task IDs, annotation for the samples in tasks is also a major concern. Almost all the available CL strategies in the medical domain require sample annotation for learning the classification or segmentation task. It is undeniable that labeling is highly costly in the medical domain, especially in the histopathology field, and hence may hinder the rapid advancement of the field. In other domains, unsupervised CL strategies have been explored; for example, Ashfahani and Pratama (2022) use only a few labeled samples to associate clusters to classes and the learning is completely unsupervised.
@@ 6 Evaluation strategy and metrics
To evaluate a CL strategy, various aspects like the performance of the model on past, current, or future episodes, resource consumption Zhang et al. (2023b), memory size Hofmanninger et al. (2020a), model size growth, execution time, etc can be explored. CL frameworks involving dynamic architecture or memory also report time and memory analysis. For example, González et al. (2022) maintains model history as mean and variance parameters of multivariate Gaussian, which dynamically grow in the presence of drift; hence, authors also provided training time analysis. If the framework does not involve dynamic architecture, primarily performance on data is reported. Further, the performance of data mainly involves evaluating and reporting the episode performance, stability, and plasticity.
In CL, there is a given sequence of episodes ($1,...,T$) to learn sequentially. Typically, we have well-defined training and test pairs as $(Tr_{i},Te_{i})$ for each $i^{th}$ episode. Upon sequential training, a train-test performance matrix $P\in\mathit{P}^{T\times T}$ is generated for the given sequence of episodes ($1,...,T$). An example matrix for $T=4$ is shown in Figure 8. The cell $P_{t,i}$ is performance on test data of $i^{th}$ episode when the model training up to $t^{th}$ episodes is complete. Thus, after training up to $t^{th}$ episodes, the performance $P_{t,i}$ for $t<i$, $t=i$, and $t>i$ correspond to past, current, and future performance, respectively.
The performance metric is chosen depending on the application. For example, IoU, dice score/coefficient, structural similarity, Hausdorff distance, average symmetric surface distance, etc, are popular choices for segmentation, whereas accuracy, recall, F1-score, and AUC are usually computed for classification applications. Once the performance metric is chosen, we can compute metrics for measuring stability and plasticity on top of it. Various metrics have been proposed to quantify stability (Table 11) and plasticity (Table 12); however, all the metrics are derived from the same train-test performance matrix (Figure 8) to measure the amount of forgetting or forward transfer.
@@ 6.1 Stability: Backward transfer
It refers to measuring the catastrophic forgetting of a model upon learning new episodes. A model is regarded as stable if it gives a similar performance for a particular episode over time. However, upon learning the new episode naively, it is possible to forget the past learning and hence disrupt the performance of the previous episode. In contrast, a carefully designed CL strategy will cause less forgetting or improvement on past episodes upon learning the new episodes. A backward transfer (BWT) metric is used to measure the amount of forgetting. In other words, it is a way to measure stability, i.e., how well the model would retain the previously acquired knowledge to prevent catastrophic forgetting.
Lopez-Paz and Ranzato (2017) define BWT as the influence of new learning $t_{i}$ task on a previous task. When learning a task $t_{i}$ increases the performance on a future task, it is $+ve$ BWT, and if it causes deterioration, it results in $-ve$ BWT. Large $-ve$ BWT is also termed as catastrophic forgetting. The BWT metric given by Lopez-Paz and Ranzato (2017) is very frequently adopted Srivastava et al. (2021); Hofmanninger et al. (2020a) and it is computed after finishing all the episodes as shown in Eq.(A) in Table 11. Average forgetting Derakhshani et al. (2021) is another popularly used metric Derakhshani et al. (2022); Chen and Tang (2022). It is computed as the average of the difference in highest performance, and the final performance reached after training on all episodes is finished (Eq.(B) in Table 11). Chee et al. (2023) compute BWT after learning each subsequent task ($t^{th}$) rather than computing it at the last episode (Eq.(B) in Table 11) as shown through Eq.(C) in Table 11. González et al. (2022) compute BWT as the change in performance after training with each subsequent task averaged over the number of tasks (Eq.(D) in Table 11). Note that BWT is not defined for the last task. Another BWT metric contribution is by Özgün et al. (2020) as shown via Eq.(E) in Table 11.
@@ 6.2 Plasticity: Forward transfer
The plasticity of a model is reflected in its capability to accommodate more and more knowledge or exploit the knowledge learned so far to learn the new. In literature, this is frequently referred to as forward transfer (FWT). It is also said that FWT measures the “zero-shot" learning capability of a model Lopez-Paz and Ranzato (2017). In contrast to the usage of BWT, not all works report FWT. Certain approaches might enforce rigorous regularization on the model to get a comparatively reduced level of forgetting (high $+ve$ BWT); however, this results in poor performance on new tasks. Therefore, reporting FWT along with BWT is equally important for unbiased evaluation of the CL model.
The very popular FWT metric Srivastava et al. (2021); Hofmanninger et al. (2020a) shown via Eq.(F) in Table 12 is proposed by Lopez-Paz and Ranzato (2017) that quantifies FWT as the impact of learning a task $t_{i}$ on a future task. It is the average of the difference in the performance on $t_{i}$ task before learning it and the performance given by a model with random weights ($\bar{b}_{i}$). Note that FWT is not defined for the first task. Further, González et al. (2022) computes FWT as the change in performance in each stage before and up to $Tr_{i}$, averaged over the number of tasks (Eq.(G) in Table 12). Another metric for FWT was proposed by Shu et al. (2022) as shown via Eq.(H) in Table 12. Özgün et al. (2020) defines two separate forward transfer metrics, one for unseen (average of elements above diagonals of $P$) and another for seen episodes as reported via Eq.(I) and Eq.(J) in Table 12, respectively. There, $b_{i}$ is the performance by a standalone model only trained on $i^{th}$ episode. Further, they report a “transfer-learning metric" as the sum of diagonal elements in matrix $P$ as measures of plasticity, i.e., the ability to adapt to new tasks.
@@ 6.3 Episode performance and others
Thus, after processing all the episodes, we get total $T$ average accuracy values which can be plotted and visualized for different approaches Derakhshani et al. (2022); Kaustaban et al. (2022). Further, researchers Derakhshani et al. (2022) also compute mean over these values ($\frac{1}{T}\sum_{i=1}^{T}[\frac{1}{t}\sum_{i=1}^{t}P_{t,i}$]) to get a single scalar value which can be used to directly compare with other approaches. Another popular definition of average accuracy is to compute the above-mentioned metric only after finishing all the episodes Lopez-Paz and Ranzato (2017); Chen and Tang (2022); Kaustaban et al. (2022) as below.
@@
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation metric Dataset
Ravishankar et al. (2019) Ultrasound, X-ray Cardiac view, Pneumothorax DIS, TIS, IIS Rehearsal + architectural CNN Supervised Accuracy, BWT Subset of ChestXRay (Wang et al. 2017)
Hofmanninger et al. (2020a) CT Chest DIS Rehearsal Res-Net50 (He et al. 2016) Supervised Accuracy, BWT, FWT Private
Lenga et al. (2020) X-ray Chest DIS Comparitive: EWC, LWF DenseNet121 (Huang et al. 2017) Supervised AUC, BWT, FWT ChestX-ray14, MIMIC-CXR
Li et al. (2020) Dermoscopic Skin disease CIS Dual-distillation+rehearsal ResNet18, AlexNet, VGG19 Supervised Accuracy Skin8, Skin40 from ISIC2019 (Tschandl et al. 2018), CIFAR100
@@
Srivastava et al. (2021) X-ray Chest DIS Rehearsal ResNet101 Supervised AUC, FWT, BWT NIH Chest-X-rays14 (McDermott et al. 2020), PadChest (Bustos et al. 2020), CheXpert (Wang et al. 2017)
Bayasi et al. (2021) Dermoscopic Skin lesion DIS Architectural (fixed but partitioned network) ResNet-50 Supervised Accuracy HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
Morgado et al. (2021) Dermoscopic Dermatological imaging modality DIS Regularization, rehearsal VGG-16, MobileNetV2 Supervised Accuracy, precision, recall, F1-score, FWT, BWT Private
@@
Shu et al. (2022) Fundus images Eyes DIS Rehearsal + regularization Resnet-18, 2-layer MLP (Saha et al. 2021) Supervised F1-score, BWT, FWT ODIR, RIADD, REFUGE, iSee (Fang et al. 2020), EyeQ (Fu et al. 2019), PMNIST (Rupesh Kumar et al. 2013)
Akundi and Sivaswamy (2022) X-ray Chest CIS Regularization DenseNet121 Supervised AUC, BWT CheXpert
Kaustaban et al. (2022) Histopathology Tumor (breast cancer, colorectal cancer) IIS, DIS, CIS, TIS Regularization, rehearsal (comparative :EWC, LwF, AGEM, CoPE, iCaRL) ResNet-18 Supervised Accuracy, FWT, BWT CRC (Kather et al. 2019), PatchCam (Bejnordi et al. 2017)
Chen and Tang (2022) Histopathology Breast cancer CIS Regularization Alexnet, Densenet201, Resnet152 Supervised Accuracy, forgetting BreakHis
Derakhshani et al. (2022) Multiple Multiple disease TIS, CIS, cross-DIS Regularization, rehearsal (comparative: EWC, MAS, LwF, iCaRL, EEIL) ResNet18 Supervised Accuracy, forgetting MedMNIST (Yang et al. 2023a)
Chee et al. (2023) Histopathology, retinal, skin lesions Colorectal cancer, eye, skin CIS Rehearsal + architecture + regularization ResNet18 Supervised Accuracy, forgetting CCH5000 (Kather et al. 2016), EyePACS (kag ), HAM10000 (Tschandl et al. 2018)
Zhang et al. (2023c) Multiple Disease (multiple organ source) CIS Rehearsal + architectural ResNet18 Supervised MCR Skin8, Path16, CIFAR100
Bándi et al. (2023) Histopathology Cancer (multi-organ: breast, colon, head-neck) DIS Comparative: (Architecture: PackNet) (Regularization: EWC, GEM) DenseNet Supervised FROC, ROC, Cohen’s Kappa CAMELYON16, CAMELYON17, private for colon and head-neck
Bai et al. (2023) VQA Surgical visual-question localized-answering CIS Regularization VisualBERT Supervised Accuracy, IoU EndoVis18, EndoVis17, M2CAI
Sadafi et al. (2023) Microscopic WBC DIS, CIS, DIS + CIS Rehearsal + regularization ResNeXt-50 Supervised Accuracy Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Yang et al. (2023b) Multiple Disease classification DIS+CIS Rehearsal ResNet18 Few-shot Accuracy, performance dropping rate MedMNIST (Yang et al. 2023a) (PathMNIST, DermaMNIST, OrganAMNIST, RetinaMNIST, BreastMNIST, BloodMNIST)
@@
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation metric Dataset
@@
Karani et al. (2018) MRI Brain DIS Architecture U-Net Supervised DSC HCP (Van Essen et al. 2013), ADNI (ADNT ), ABIDE (Di Martino et al. 2014) and IXI (IXI )
@@
Baweja et al. (2018) MRI Normal brain structures, white matter lesions TIS Regularization: EWC DeepMedic 3D convnet (Kamnitsas et al. 2017) Supervised DSC coefficient UK Biobank (Miller et al. 2016)
McClure et al. (2018) SMRI Axial and sagittal brain DIS Regularization: DWC MeshNet (Fedorov et al. 2017) (CNN-based) Supervised DSC HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017)
Ozdemir et al. (2018) MRI Humerus and scapula CIS Regularization (LwF) + rehearsal U-Net Supervised DSC, SSD Private
@@
Ozdemir and Goksel (2019) MRI Knee CIS Regularization U-Net Supervised DSC, MSD SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
van Garderen et al. (2019) 4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR Brain tumor DIS Regularization: EWC 3D U-Net CNN Supervised Dice score BraTS 2018 (Bakas et al. 2017; Menze et al. 2014), private
@@
Özgün et al. (2020) MRI Brain DIS Regularization: MAS QuickNAT (Roy et al. 2019) Supervised CL-DSC, REM, BWT+, TL, FWT CANDI (Kennedy et al. 2012), MALC (Asman and Landman 2013), and ADNI (Jack Jr et al. 2008)
Memmel et al. (2021) MRI Hippocampal DIS U-Net Supervised IoU, DSC (Simpson et al. ), (Kulaga-Yoskovitz et al. 2015), (Boccardi et al. 2015)
Perkonigg et al. (2021) MRI Brain DIS Rehearsal 3D-ModelGenesis (Zhou et al. 2021b) Supervised MAE, FWT, BWT IXI (IXI ), OASIS3 (LaMontagne et al. 2019)
@@
Ranem et al. (2022) MRI Hippocampus DIS Comparative (rehearsal, regularization, hybrid) ViT U-Net Supervised DSC, FWT, BWT Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
@@
Liu et al. (2022) CT Liver, spleen, pancreas, right kidney left kidney CIS Rehearsal + regularization nnUNet Supervised DC, HD95 (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
@@
Karthik et al. (2022) MRI Multiple sclerosis lesions (brain) DIS Rehearsal 3D U-Net Supervised DSC, BWT (Kerbrat et al. 2020)
@@
González et al. (2022) MRI Hippocampus DIS Architectural nnUNet (Isensee et al. 2021) Supervised DSC, FWT, BWT HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019)
Li et al. (2022) CMR Cardiac DIS Rehearsal U-Net Supervised DSC, HD95, BWT, FWT M&Ms (Campello et al. 2021)